In the 2016 Olympics in Rio, after the 50 m freestyle competition, a problem with the pool was found. In lane 1 there was a gentle 1.2 cm/s current flowing in the direction that the swimmers were going, while in lane 8 there was a current of the same speed but directed opposite to the swimmers' direction. Suppose a swimmer could swim the 50.0 m in 25.0 s in the absence of any current.
How would the time it took the swimmer to swim 50.0 m change in lane 1?
1
Expert's answer
2019-10-02T09:37:39-0400
"t'=\frac{d}{V+u}"

"V=\frac{d}{t}"

Thus, the time decreases by


"t-t'=t-\frac{d}{\frac{d}{t}+u}=t\left(1-\frac{1}{1+\frac{ut}{d}}\right )"


"t-t'=25\left(1-\frac{1}{1+\frac{(0.012)(25)}{50}}\right )=0.15\ s"